OpenStudy (anonymous):

Use substitution, addition/subtraction, or linear combinations to solve the linear system. 10r-9s=18 6s+2r=1

hero (hero):

Hi

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

any help?

hero (hero):

Yes, I will solve it

hero (hero):

Give me a minute

OpenStudy (anonymous):

ok

hero (hero):

r = 3/2, s = -1/3

OpenStudy (anonymous):

easy